OpenStudy (anonymous):

Finding the Percentage 7.4% of $694.56 -$93.86 -$51.40 -$938.59 8 ¼ % of $2,458.20 -$297.96 -$2,028.02 -$202.80

OpenStudy (jdoe0001):

\(\large {{\color{red}{ a}}\%\ of \ b=\cfrac{{\color{red}{ a}}}{100}\times b\\ \quad \\ {\color{red}{ 7.4}}\%\ of \ 694.56=\cfrac{{\color{red}{ 7.4}}}{100}\times 694.56\\ \quad \\ {\color{red}{ 8\frac{1}{4}}}\%\ of \ 2,458.20=\cfrac{{\color{red}{ 8\frac{1}{4}}}}{100}\times 2,458.20}\)